OpenStudy (anonymous):

If f(x) varies directly with x2, and f(x) = 75 when x = 5, find the value of f(4). (Points : 1) 100 50 48 25

Nnesha (nnesha):

\[\huge\rm y =kx^2\] equation for direct variation substitute x and y values then solve for k

Nnesha (nnesha):

F(X) same as y

OpenStudy (here_to_help15):

f(x)= a*x^2 and f(5)=75 75=a*5^2 \[\rightarrow f(x) =3x^2 \rightarrow f(4)= 3*4^2=3*16=48\]

OpenStudy (here_to_help15):

Oops a=75/25=3

OpenStudy (here_to_help15):

Question did i give a direct answer?

OpenStudy (here_to_help15):

@eliassaab @Gabylovesyou @Compassionate @bohotness @nincompoop @Phebe @phi @StudyGurl14 @undeadknight26

OpenStudy (here_to_help15):

@Nnesha said i have gave a direct mod, did i give a direct answer 0.0 ?

OpenStudy (here_to_help15):

direct answer* haha

Nnesha (nnesha):

ohh i want to keep that private if u really want to create a issue than let's tag mod instead of regular users

Nnesha (nnesha):

then**

OpenStudy (studygurl14):

Yes, you did. It is good that you explained your answer, but you're supposed to guide the user to the answer and try to get them to do some of the work themselves.

OpenStudy (here_to_help15):

haha sorry @Nnesha i didnt mean for no negativity between each other just was wondering

OpenStudy (here_to_help15):

:(

Nnesha (nnesha):

i have different ways to help other users i know this user is learning about "direct variation" last question was almost same like this one if we give her a chance to do it then she don't have to post same type of question again and again and as a user of os WE DON'T WANT THEM TO BE DEPENDENT ON OTHER USERS our purpose to make them independent so they can work their own instead of waiting someone for help or then they can be able to help other users

OpenStudy (studygurl14):

^True. I've seen users post like 5 questions in a row that are almost identical.